OpenStudy (anonymous):

Find the limit. Use l'Hospital's Rule where appropriate. Lim x->0 [ sqrt(1+3x) - sqrt(1-5x) ] / x

OpenStudy (anonymous):

Find the Derivative of Numerator and Denominator...Seperately

OpenStudy (anonymous):

Yeah I tried but the denominator is still =0

OpenStudy (anonymous):

d (x)/dx = 1

OpenStudy (anonymous):

so denominator is 1

OpenStudy (anonymous):

You used the quotient rule?

OpenStudy (anonymous):

No...) Differentiate Seperately..

OpenStudy (anonymous):

sqrt(1+3x) = (1 + 3x) ^1/2

OpenStudy (anonymous):

if you do difference separately, then you need to divide both by x which will need the quotient rule as well. I entered 1 as my answer but it's wrong :(

OpenStudy (anonymous):

In L'Hopital's rule Lt {f(x)/g(x)}=Lt {f'(x)/g'(x)} So you'll need to differentiate both numerator n denominator separately!

OpenStudy (anonymous):

Okay I got my answer. -1?

OpenStudy (anonymous):

Answer should be 4

OpenStudy (anonymous):